i keep doing this matrices
2x-y+9x=-8
-x-3y+4z=-15
5x+2y-z=17
and i get a fraction at the end can you help me?

Answer by richwmiller(17219)

2x-y+9x=-8
-x-3y+4z=-15
5x+2y-z=17
Using Gauss Jordan
2,-1,9,-8
-1,-3,4,-15
5,2,-1,17
divide row 1 by 2
1,-1/2,9/2,-4
-1,-3,4,-15
5,2,-1,17
add down (1) *row 1 to row 2
1,-1/2,9/2,-4
0,-7/2,17/2,-19
5,2,-1,17
add down (-5) *row 1 to row 3
1,-1/2,9/2,-4
0,-7/2,17/2,-19
0,9/2,-47/2,37
divide row 2 by -7/2
1,-1/2,9/2,-4
0,1,34/-14,-38/-7
0,9/2,-47/2,37
add down (-9/2) *row 2 to row 3
1,-1/2,9/2,-4
0,1,17/-7,38/7
0,0,352/-28,176/14
divide row 3 by 88/-7
1,-1/2,9/2,-4
0,1,17/-7,38/7
0,0,1,-1
We now have the value for the last variable.
We will work our way up and get the other solutions.
add up (-17/-7) *row 3 to row 2
1,-1/2,9/2,-4
0,1,0,3
0,0,1,-1
add up (-9/2) *row 3 to row 1
1,-2/4,0,1/2
0,1,0,3
0,0,1,-1
add up (1/2) *row 2 to row 1
1,0,0,2
0,1,0,3
0,0,1,-1
final
1,0,0,2
0,1,0,3
0,0,1,-1
1,0,0,2
0,1,0,3
0,0,1,-1
"2","3","-1"
(2,3,-1)